OpenStudy (venesa):

prove that cos(a+b).cos(a-b) = 1-(sin^a+sin^b)

OpenStudy (anonymous):

^ does this mean to the power ?

OpenStudy (venesa):

it is the square

OpenStudy (anonymous):

cos(a+b).cos(a-b)=(cos2a+cos2b)/2 =1/2*[1-2sin^2a+1-2sin^2b] =1/2*2[1-(sin^2a+cos^2b)] =1-(sin^2a+cos^2b)

OpenStudy (anonymous):

did you get the previous solution?

OpenStudy (venesa):

* does it mean multiplication ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you could work this way too cos(A+B).cos(A-B) =(cosA cos B - sinAsinB)(cosA cos B +sinAsinB) and then multiply the terms

OpenStudy (venesa):

@RED123url : i can't answer the questions :(